OpenStudy (anonymous):

the equation of a parabola with vertex (0, 0) and directrix x = -3

OpenStudy (jdoe0001):

so the parabola looks like now your center is at (0, 0) use the focus standard form for the parabola with a focus distance at "p" \(\huge (x-h)^2=4p(y-k)\)

OpenStudy (jdoe0001):

the focus is "3", because in a parabola the directrix and focus are at EXACTLY the same distance from the vertex

OpenStudy (jdoe0001):

if the directrix is at -3, the focus will be on the other end, +3

OpenStudy (anonymous):

So then it'd be (x - h)^2 = 4(3) (y - k) ?

OpenStudy (jdoe0001):

yes

jimthompson5910 (jim_thompson5910):

jdoe000, the directrix is a vertical line, not a horizontal line

OpenStudy (jdoe0001):

since your vertex, no center my bad, your vertex is (0, 0), that'd be simple you can then solve if you wish for "y"

jimthompson5910 (jim_thompson5910):

the axis of symmetry and the directrix are not the same line

OpenStudy (jdoe0001):

ohh shoot, yeah, hold the mayo

OpenStudy (jdoe0001):

@jim_thompson5910 is correct,, the directrix is x = -3 so is really on the other direction